# Quantised helicity in optical media

### v1

@@ IV Diagonal form of the helicity
$\hskip 56.9055pt+i\left(\hat{c}_{\lambda}(\mathbf{k})\hat{b}^{\dagger}_{\lambda}(\mathbf{k})-\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\hat{b}_{\lambda}(\mathbf{k})\right).$
This can further be rewritten in terms of polaritons using Eqns. (SI 6), where we should note that we need to normal order the operators with respect to the polariton vacuum. This yields
$:\hat{h}_{\text{total}}:=\sum_{\lambda=\pm}$ $\int d^{3}k\;\sum_{ij}\bigg{(}\lambda\left(\alpha_{i}\alpha_{j}^{*}+\widetilde{\alpha}_{i}\widetilde{\alpha}_{j}^{*}\right)$
@@
$+2\text{Im}\left[\beta_{i}^{*}\gamma_{j}+\widetilde{\gamma}_{i}^{*}\widetilde{\beta}_{j}\right]\bigg{)}\hat{p}^{\dagger}_{i}(\mathbf{k},\lambda)\hat{p}_{j}(\mathbf{k},\lambda)$
+$+\bigg{\{}\left(2\text{Im}\left[\widetilde{\gamma}_{i}\beta_{j}-\gamma_{i}\widetilde{\beta}_{j}\right]-\lambda\alpha_{i}\widetilde{\alpha}_{j}\right)$
[…]
@@
We can simplify this by using Eq. (III) to express $\beta_{i}$, $\widetilde{\beta}_{i}$, $\gamma_{i}$ and $\widetilde{\gamma}_{i}$ in terms of $\alpha_{i}$ and $\widetilde{\alpha}_{i}$. We may then use the fact that $\widetilde{\alpha}_{u/d}=0$ and $\alpha_{p}=0$, as well as the explicit solutions for $\alpha_{u/d}$ and $\widetilde{\alpha}_{p}$ given in Eq. (SI 12) and Eq. (SI 13), to show that the $i\neq j$ terms vanish.
@@
In order to show this cancellation, we also must make use of the fact that when $i\neq j$, the frequencies $\omega_{i}$ and $\omega_{j}$ are related to one another, as both are two distinct roots of the same dispersion relation. The relevant dispersion relations Eqns. (SI 10) and  (SI 11) are cubics in $\omega$, so sums and products of the two roots which appear can be expressed in terms of the third root. Finally, using Eq. (SI 12) and Eq. (SI 13) again, we find Eq. (Quantised helicity in optical media) from the main text.

### v2

@@ Appendix D Diagonal form of the helicity
$\hskip 28.45274pt+\frac{i}{2}\Bigg(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega_{E}}}{\tilde{\omega}_{M}}}+\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega_{M}}}{\tilde{\omega_{E}}}}\Bigg)\left(\hat{c}_{\lambda}(\mathbf{k})\hat{b}^{\dagger}_{\lambda}(\mathbf{k})-\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\hat{b}_{\lambda}(\mathbf{k})\right)$
@@
$\hskip 28.45274pt-\frac{i}{2}\Bigg(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega_{E}}}{\tilde{\omega}_{M}}}-\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega_{M}}}{\tilde{\omega_{E}}}}\Bigg)\left(\hat{b}_{\lambda}(-\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})-\hat{b}^{\dagger}_{\lambda}(-\mathbf{k})\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\right).$ (52)
This can further be rewritten in terms of polaritons using Eqns. (44), where we should note that we need to normal order the operators with respect to the polariton vacuum. This yields
$\int{d^{3}r:\hat{h}:}=\int\,d^{3}k\sum_{\lambda}\sum_{ij}$ $\hat{p}^{\dagger}_{i}(\mathbf{k},\lambda)\hat{p}_{j}(\mathbf{k},\lambda)\Bigg(\lambda\alpha_{i}\alpha^{*}_{j}+\lambda\tilde{\alpha}^{*}_{j}\tilde{\alpha}_{i}+\frac{i}{2}\Big(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}+\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Big)(\gamma^{*}_{j}\beta_{i}+\tilde{\gamma}_{i}\tilde{\beta}_{j}^{*}-\gamma_{i}\beta_{j}^{*}-\tilde{\gamma}_{j}^{*}\tilde{\beta}_{i})$
$-\frac{i}{2}\Big(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}-\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Big)(-\tilde{\beta}_{i}\gamma_{j}^{*}-\beta_{j}^{*}\tilde{\gamma}_{i}+\beta_{i}\tilde{\gamma}_{j}^{*}+\tilde{\beta}_{j}^{*}\gamma_{i})\Bigg)$
$+\Bigg[$ $\hat{p}_{i}(-\mathbf{k},\lambda)\hat{p}_{j}(\mathbf{k},\lambda)\Big(-\frac{\lambda}{2}(\tilde{\alpha}_{i}^{*}\alpha_{j}^{*}+\alpha_{i}^{*}\tilde{\alpha}_{j}^{*})+\frac{i}{2}\Big(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}+\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Big)(-\gamma_{i}^{*}\tilde{\beta}_{j}^{*}+\tilde{\gamma}_{i}^{*}\beta_{j}^{*})$
- $\frac{i}{2}\Big(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}-\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Big)(\gamma_{i}^{*}\beta_{j}^{*}-\tilde{\gamma}_{i}^{*}\tilde{\beta}_{j}^{*})\Big)+\mathrm{h.c}\Bigg].$ (53)
In the dual-symmetric case, this simplifies to
$\begin{split}\int{d^{3}r:\hat{h}_{D.S.}:}\,&=\sum_{\lambda}\int{d^{3}k\sum_{i,j}\hat{p}_{i}^{\dagger}(\mathbf{k})\hat{p}_{j}(\mathbf{k})\Big(\lambda(\alpha_{i}\alpha_{j}+\tilde{\alpha}_{i}\tilde{\alpha}_{j})+i(\beta_{i}\gamma_{j}-\tilde{\gamma}_{i}\tilde{\beta}_{j}+\gamma_{i}\beta_{j}-\tilde{\beta}_{i}\tilde{\gamma}_{j})\Big)}\\ &+\hat{p}^{\dagger}_{i}(\mathbf{k})\hat{p}^{\dagger}_{j}(-\mathbf{k})\Big(-\lambda\alpha_{i}\tilde{\alpha}_{j}+i(\gamma_{i}\tilde{\beta}_{j}-\tilde{\gamma}_{i}\beta_{j})\Big)\\ &+\hat{p}_{i}(\mathbf{k})\hat{p}_{j}(-\mathbf{k})\Big(-\lambda\tilde{\alpha}_{i}\alpha_{j}+i(\gamma_{i}\tilde{\beta}_{j}-\tilde{\gamma}_{i}\beta_{j})\Big),\end{split}$ (54)
where we have made use of the fact that we are free to choose the phases of the coefficients such that $\alpha$ and $\tilde{\alpha}$ are real. The expressions  (C) then immediately imply that $\beta$ and $\tilde{\beta}$ are purely imaginary, while $\gamma$ and $\tilde{\gamma}$ are purely real.
@@
We can simplify this further by using Eq. (C) to express $\beta_{i}$, $\widetilde{\beta}_{i}$, $\gamma_{i}$ and $\widetilde{\gamma}_{i}$ in terms of $\alpha_{i}$ and $\widetilde{\alpha}_{i}$. We may then use the fact that $\widetilde{\alpha}_{u/d}=0$ and $\alpha_{p}=0$, as well as the explicit solutions for $\alpha_{u/d}$ and $\widetilde{\alpha}_{p}$ given in Eq. (50) and Eq. (51), to show that the $i\neq j$ terms vanish.
@@
In order to show this cancellation, we also must make use of the fact that when $i\neq j$, the frequencies $\omega_{i}$ and $\omega_{j}$ are related to one another, as both are two distinct roots of the same dispersion relation. The relevant dispersion relations Eqns. (48) and  (49) are cubics in $\omega$, so sums and products of the two roots which appear can be expressed in terms of the third root. Finally, using Eq. (50) and Eq. (51) again, we find Eq. (29) from the main text.
Appendix E Diagonalisation in non-dual-symmetric media
Here we present explicit expressions for the polariton operators in a non-dual-symmetric medium. It is possible to exactly diagonalise the Hamiltonian (25), without making any assumptions about $\omega_{E}$, $\omega_{M}$, $\alpha$ or $\beta$. If we do so, we obtain the polariton operators
$\hat{p}_{\lambda_{i}}(\mathbf{k})=\alpha_{i}a_{\lambda}(\mathbf{k})+\tilde{\alpha}_{i}a_{\lambda}^{\dagger}(-\mathbf{k})+\beta_{i}b_{\lambda}(\mathbf{k})+\tilde{\beta}_{i}b_{\lambda}^{\dagger}(-\mathbf{k})+\gamma_{i}c_{\lambda}(\mathbf{k})+\tilde{\gamma}_{i}c_{\lambda}^{\dagger}(-\mathbf{k}),$ (55)
with coefficients
$|\alpha_{i}|^{2}$ $=\Bigg(1-A^{2}+\frac{\alpha^{2}k(1+A)^{2}\omega_{i}}{(\omega_{i}-\tilde{\omega}_{E})^{2}(\omega_{i}+\tilde{\omega}_{E})^{2}}+\frac{\beta^{2}k(A-1)^{2}\omega_{i}}{(\omega_{i}-\tilde{\omega}_{M})^{2}(\omega_{i}+\tilde{\omega}_{M})^{2}}\Bigg)^{-1},$ (56)
$\tilde{\alpha}_{i}$ $=A\alpha_{i},$ (57)
$\beta_{i}$ $=\frac{i\alpha}{2(\omega_{i}-\tilde{\omega}_{E})}\sqrt{\frac{k}{\tilde{\omega}_{E}}}(1+A)\alpha_{i},$ (58)
$\gamma_{i}$ $=\frac{\lambda\beta}{2(\omega_{i}-\tilde{\omega}_{M})}\sqrt{\frac{k}{\tilde{\omega}_{M}}}(A-1)\alpha_{i},$ (59)
$\tilde{\beta}_{i}$ $=\frac{i\alpha}{2(\omega_{i}+\tilde{\omega}_{E})}\sqrt{\frac{k}{\tilde{\omega}_{E}}}(1+A)\alpha_{i},$ (60)
$\tilde{\gamma}_{i}$ $=\frac{\lambda\beta}{2(\omega_{i}+\tilde{\omega}_{M})}\sqrt{\frac{k}{\tilde{\omega}_{M}}}(A-1)\alpha_{i}$ (61)
with
$A\equiv\Bigg(\frac{\beta^{2}k}{2(\omega_{i}^{2}-\tilde{\omega}_{M}^{2})}-\frac{\alpha^{2}k}{2(\omega_{i}^{2}-\tilde{\omega}_{E}^{2})}\Bigg)\Bigg(\omega_{i}+k+\frac{\alpha^{2}k}{2(\omega_{i}^{2}-\tilde{\omega}_{E}^{2})}+\frac{\beta^{2}k}{2(\omega_{i}^{2}-\tilde{\omega}_{M}^{2})}\Bigg)^{-1}.$ (62)
The phases of the coefficients $\alpha_{i}$ may be chosen freely. In contrast with the dual-symmetric case presented in Appendix C, we no longer have that $\alpha_{i}\neq 0\implies\tilde{\alpha}_{i}=0$, nor do we generally have one negatively refracting branch.
Appendix F Eigenvectors of the Duality Transform
In this appendix, we discuss the field vectors which are eigenvectors of the duality transformation. We have seen that the helicity per circularly polarised polariton is not $\pm 1$ in non-dual-symmetric media. One way to understand this is by considering, classically, the field vectors which are eigenvectors of the duality transform and the circumstances under which these coincide with the fields of a circularly polarised plane wave in a linear magnetodielectric.
Firstly, we return to the fact that the duality transform (II) can be viewed as a rotation between electric and magnetic quantities. This is most apparent when written in matrix form,
$\left(\begin{array}[]{cc}\mathbf{A}^{\prime}\\ \mathbf{C}^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{array}\right)\left(\begin{array}[]{cc}\mathbf{A}\\ \mathbf{C}\end{array}\right),$ (69)
$\left(\begin{array}[]{cc}\mathbf{P}^{\prime}\\ \mathbf{M}^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{array}\right)\left(\begin{array}[]{cc}\mathbf{P}\\ \mathbf{M}\end{array}\right).$ (76)
The eigenvectors [^2] There is also a $z$-directional eigenvector of the duality transform, which is not relevant here as we are discussing transverse fields. of the duality rotation (76) can be written as
$\mathbf{e}^{\text{duality}}_{\lambda}=\frac{1}{2}\left(\begin{array}[]{cc}\mathbf{e}_{x}+i\lambda\mathbf{e}_{y}\\ -i\lambda\mathbf{e}_{x}+\mathbf{e}_{y}\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}\mathbf{e}_{\lambda}\\ -i\lambda\mathbf{e}_{\lambda}\end{array}\right),$ (81)
with eigenvalues $e^{-i\lambda\theta}$. We can then connect this to rotations in physical space by considering a circularly polarised plane wave propagating in the $z$-direction. Let us first return to free-space, where an electric field $\mathbf{E}=E_{0}\mathbf{e}_{\lambda}$ may be described by field vectors $\mathbf{A}_{0}\propto\mathbf{e}_{\lambda}$ and $\mathbf{C}_{0}=-i\lambda\mathbf{A}_{0}\propto-i\lambda\mathbf{e}_{\lambda}$, with $\lambda=\pm 1$ referring to the two circular polarisations. The vector $(\mathbf{A}_{0},\mathbf{C}_{0})^{T}$ is then clearly proportional to the duality transform eigenvector (81). These are, as can be seen, closely connected to the eigenvectors of physical rotation about the $z$-direction, which are the circular polarisation vectors $\mathbf{e}_{\pm}=\mathbf{e}_{x}\pm i\mathbf{e}_{y}$, also with eigenvalues $e^{\pm i\theta}$.
In a dual-symmetric medium, the $\mathbf{A}$ and $\mathbf{C}$ fields are still related as in free-space and induce polarisations and magnetisations $\mathbf{P}_{0}\propto\mathbf{e}_{\lambda}$ and $\mathbf{M}_{0}=-i\lambda\mathbf{P}_{0}\propto-i\lambda\mathbf{e}_{\lambda}$. Because the polarisation and magnetisation have equal amplitude, the vectors $(\mathbf{P}_{0},\mathbf{M}_{0})$ of this form are also eigenvectors of the $(\mathbf{P},\mathbf{M})$ transform. In this sense, the helicity quantifies the twist associated with the rotation implied by the circular polarisation vectors.
However, this is not the case in general. In most optical media, the potentials $\mathbf{A}$ and $\mathbf{C}$ for a circularly polarised wave are no longer related as simply. These then also give rise to polarisation and magnetisation fields that are not related by $\mathbf{M}=-i\lambda\mathbf{P}$. As an example of this, let us consider a medium with permittivity $\varepsilon$, permeability $\mu$ and susceptibilities $\chi_{e}$ and $\chi_{m}$. The potentials are in this case related by $\mathbf{C}=-i\lambda\sqrt{\varepsilon/\mu}\,\mathbf{A}$. Also, the polarisation and magnetisation are instead related by $\mathbf{M}=-i\lambda(\chi_{m}/\chi_{e})\sqrt{\mu/\varepsilon}\,\mathbf{P}$. This does not form an eigenvector of the duality transform, so the circular polarisations are no longer the states of unit helicity. Furthermore, in a dispersive medium where the susceptibilities are functions of frequency, we can see that the scaling between $\mathbf{M}$ and $\mathbf{P}$ will generally be $k$-dependent. This is what causes the helicity of a circularly polarised wave to depend on $k$, as those vectors can always be decomposed in terms of $\mathbf{e}^{\text{duality}}_{\pm}$ with $k$-dependent amplitudes. [^3] This is a change of basis, for which the $z$-directional eigenvector is required in the definition. It does not enter in the decomposition, however.
Another way to put this is that, for a plane wave in a dual-symmetric medium, the rotation (II) corresponds to a physical rotation of both the fields and the dipoles about the direction of propagation. In a general medium, this is not the case. It is possible to scale the transform as in (II) to retain the connection between the transformation and rotations about $\mathbf{k}$ in some cases, but this is not possible in general. We choose in this work to use the unscaled transformation in all cases, and to interpret the differences in behaviour between helicity in media and in free space as a manifestation of the fact that the matter and light behave in different ways. We emphasise that, although we have considered linear magnetodielectrics in this discussion, the application of (II) and its associated helicity (9) does not depend on any particular constitutive relation connecting the light and matter fields.